# HATS-38 b and WASP-139 b join a growing group of eccentric hot Neptunes on polar orbits [^1] Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programs ID 111.24VT.001, 111.24VT.002, and 112.25W1.001

### v1

@@ 6.3 Possible Origin of the Polar Orbits
We discuss some of the proposed explanations for the origin of the polar orbits of hot Neptunes and whether these are able to explain not only the stellar obliquities of HATS-38 b and WASP-139 b, but also their residual eccentricities and constraints on long-period companions. We argue that primordial misalignments and in situ excitation are ruled out by these observational constraints, while high-eccentricity migration from $\gtrsim 2$ au remains as a possible formation pathway.
@@
Highly misaligned orbits can be produced through a primordial tilt of the protoplanetary disk by gravitational torques from a binary stellar companion (e.g., Batygin, 2012; Lai, 2014; Zanazzi & Lai, 2018). This scenario seems possible given that both stars appear to have distant M-dwarf companions at projected separations of $\sim 2100$ au and $\sim 6400$ au (Section 6.2). However, at these large separations a precession cycle of the disk is long [^7] For reference, a protoplanetary disk extending out to 200 au would complete a precession cycle in $\sim 15$ Myr due to a M-dwarf companion with semi-major axis of 2,000 au (e.g., Equation (12) in Zanazzi & Lai 2018). compared to the disk dispersal timescale to resonantly excite the stellar obliquities (Zanazzi & Lai, 2018). Alternatively, the stellar obliquity may be excited by the disk through magnetic disk torquing in a young accreting star (Lai et al., 2011). Although these may still be plausible processes to misalign the parent disks of HATS-38 b and WASP-139 b to the equator of their host stars, they are unable to explain the observed eccentricities. Therefore, we disfavor primordial misalignments as an explanation.
@@
For HATS-38 b and WASP-139 b, we obtain $\theta^{2}\simeq 0.02$ and $\theta^{2}\simeq 0.04$, respectively, implying in both cases that scattering leads mainly to planetary collisions, not ejections. Thus, the excitation of eccentricities and inclinations becomes unlikely at the planet’s orbital separation (Petrovich et al., 2014).
@@
For $\eta_{\rm GR}>3$ eccentricity excitation is quenched. In Figure 5, we show that the allowed regions for HATS-38 b and WASP-139 b (pink dot-dashed lines) are largely excluded by long-term RV measurements.
@@
For $\eta_{\star}>1$ a distant planet "c" is unable to drive the growth of the stellar obliquity. In both panels of Figure 5 we plot $\eta_{\star}=1$ assuming $J_{2}=10^{-7}$ and the observed stellar masses and radii, and find that we can rule out the presence of a planetary companion in most of the regions where the necessary condition is satisfied.
@@
We have ruled out various mechanisms that can explain the excited orbital states of the Neptunes at their current orbital distances. We turn to a final possibility, namely excitation of their stellar obliquities and eccentricities at wider orbital separations, followed by inward tidal migration.
@@
If no companions are found, we are left with the wide-orbit binary companions as the driver of ZLK migration (e.g., Fabrycky & Tremaine 2007; Naoz et al. 2012; Petrovich 2015), though migration should have started from an even wider orbit to not be quenched by GR ($\eta_{\rm GR}<1/2$ for $a_{\rm b,0}\gtrsim 9$ au with $m_{c}=0.1$ $M_{\odot}$ and $a_{c}=2000$ au, see Section 6.2). Additionally, the migrating planets should lack any other planetary companions to avoid quenching the ZLK cycles, a state that could be achieved by planet-planet scattering.
@@ 7 Conclusions and Summary
In this work, we have presented new observations of the hot Neptunes HATS-38 b and WASP-139 b. We have jointly analyzed ESPRESSO observations of the RM effect, produced by the transiting planets, with photometry and out-of-transit radial velocities using the publicly available code ironman. We have concluded that:
  • Both planets have nearly polar and eccentric orbits. We have measured a sky-projected obliquity $\lambda=-108_{-16}^{+11}$ and $-85.6_{-4.2}^{+7.7}$ deg, and an eccentricity $e=0.112_{-0.070}^{+0.072}$ and $0.103_{-0.041}^{+0.050}$ for HATS-38 b and WASP-139 b, respectively.
@@
  • Both planets join a growing population of isolated low-density Neptunes in polar and eccentric orbits, which includes HAT-P-11 b (Sanchis-Ojeda & Winn, 2011), GJ 436 b (Bourrier et al., 2018, 2022), WASP-107 b (Dai & Winn, 2017; Rubenzahl et al., 2021), and GJ 3470 b (Stefansson et al., 2022).
  • A possible explanation for the polar and eccentric architectures of these systems is high-eccentricity migration produced by secular interactions from yet undetected outer planetary companions or the observed wide-orbit stars. In contrast, we disfavor in-situ excitation of inclinations and eccentricities, as this would require companions at orbital distances we have ruled out with out-of-transit RVs.
@@ Acknowledgments
We would like to thank Coel Hellier for the TRAPPIST and EulerCAM light curves of WASP-139. We also thank Jiayin Dong for her help with the implementation of the Bayesian framework to derive the true 3D obliquity distributions. J.I.E.R and C.P. acknowledge support from ANID BASAL project FB210003. J.I.E.R. acknowledges support from ANID Doctorado Nacional grant 2021-21212378. C.P. acknowledges support from FONDECYT Project 1210425, CASSACA grant CCJRF2105, and ANID+REC Convocatoria Nacional subvención a la instalación en la Academia convocatoria 2020 PAI77200076. A.J. and R.B. acknowledge support from ANID – Millennium Science Initiative – ICN12_009. R.B. acknowledges support from FONDECYT Project 1241963. A.J. acknowledges support from FONDECYT project 1210718. J.P.L. acknowledges co-funding by the European Union (ERC, FIERCE, 101052347), by FCT - Fundação para a Ciência e a Tecnologia - through national funds, and by FEDER through COMPETE2020 - Programa Operacional Competitividade e Internacionalização grants UIDB/04434/2020 and UIDP/04434/2020.

### v2

@@ 6.3 Possible Origin of the Polar Orbits
We discuss some of the proposed explanations for the origin of the polar orbits of hot Neptunes and whether these are able to explain not only the stellar obliquities of HATS-38 b and WASP-139 b but also their residual eccentricities (potential eccentricity in the case of HATS-38 b) and constraints on long-period companions. We argue that high-eccentricity migration from $\gtrsim 2$ au is a possible formation pathway for both systems, while primordial misalignments can be an option for HATS-38 b if its orbit is circular. In-situ excitation of obliquities and eccentricities is ruled out by our observational constraints.
@@
Highly misaligned orbits can be produced through a primordial tilt of the protoplanetary disk by gravitational torques from a binary stellar companion (e.g., Batygin, 2012; Lai, 2014; Zanazzi & Lai, 2018). This scenario seems possible given that both stars appear to have distant M-dwarf companions at projected separations of $\sim 2100$ au and $\sim 6400$ au (Section 6.2). However, at these large separations a precession cycle of the disk is long [^7] For reference, a protoplanetary disk extending out to 200 au would complete a precession cycle in $\sim 15$ Myr due to a M-dwarf companion with a semi-major axis of 2000 au (e.g., Equation (12) in Zanazzi & Lai 2018). compared to the disk dispersal timescale to resonantly excite the stellar obliquities (Zanazzi & Lai, 2018). Alternatively, the stellar obliquity may be excited by the disk through magnetic disk torquing in a young accreting star (Lai et al., 2011). Although these may be plausible processes to misalign the parent disks of planets to the equator of their host stars, they are unable to explain the eccentric orbits. Given that we can not tell if the orbit of HATS-38 b is eccentric or not, these processes might be a likely explanation for its polar orbit only if the orbit is circular. If the orbit is eccentric, like in the case of WASP-139 b, these processes are less likely.
@@
For HATS-38 b and WASP-139 b, we obtain $\theta^{2}\sim 0.02$ and $\sim 0.04$, respectively, implying in both cases that scattering leads mainly to planetary collisions, not ejections. Thus, the excitation of eccentricities/inclinations becomes unlikely at the planet’s orbital separation (Petrovich et al., 2014).
@@
For $\eta_{\rm GR}>3$ eccentricity excitation is quenched. In Figure 5, we show that the allowed regions for HATS-38 b and WASP-139 b (yellow dot-dashed lines) are largely excluded by long-term RV measurements.
@@
For $\eta_{\star}>1$ the distant planet c is unable to drive the growth of the stellar obliquity. In both panels of Figure 5 we plot $\eta_{\star}=1$ assuming $J_{2}=10^{-7}$ and the observed stellar masses and radii, and find that we can rule out the presence of a planetary companion in most of the regions where the necessary condition is satisfied.
@@
We have ruled out various mechanisms that can explain the excited orbital states of the Neptunes at their current orbital distances, except for a primordial misalignment in the case of a circular orbit for HATS-38 b. Finally, we explore the possibility of high-eccentricity tidal migration: the excitation of the stellar obliquities and eccentricities at wider orbital separations, followed by inward tidal migration that circularizes the orbits over time.
@@
If no companions are found, we are left with the wide-orbit binary companions as the driver of ZLK migration (e.g., Fabrycky & Tremaine 2007; Naoz et al. 2012; Petrovich 2015). Though ZLK migration starting in itself does not predict an excess of polar planets, the combination of disk tilting due to the binary and subsequent ZLK migration can produce an excess of polar and retrograde hot planets (Vick et al., 2023). In this scenario, migration should have started from an even wider orbit to not be quenched by GR ($\eta_{\rm GR}<1/2$ for $a_{\rm b,0}\gtrsim 9$ au with $m_{c}=0.1$ $M_{\odot}$ and $a_{c}=2000$ au, see Section 6.2). Additionally, the migrating planets should lack any other planetary companions to avoid quenching the ZLK cycles, a state that could be achieved by planet-planet scattering.
@@ 7 Conclusions and Summary
In this work, we have presented new observations of the hot Neptunes HATS-38 b and WASP-139 b. We have jointly analyzed ESPRESSO observations of the RM effect produced by the transiting planets, with photometry and out-of-transit radial velocities using the publicly available code ironman. We have concluded that:
  • HATS-38 b has a polar and potentially eccentric orbit. We have measured a sky-projected obliquity $\lambda=-108_{-16}^{+11}$ deg for the orbit, and an eccentricity of $e=0.112_{-0.070}^{+0.072}$ that is compatible with 0 given our data. Future RV monitoring and/or eclipse observations will be needed to robustly distinguish between an eccentric and circular orbit for this planet.
  • WASP-139 b has a polar and eccentric orbit. We have measured a sky-projected obliquity $\lambda=-85.6_{-4.2}^{+7.7}$ deg, and an eccentricity $e=0.103_{-0.041}^{+0.050}$.
@@
  • Both planets join a growing population of isolated low-density Neptunes in polar orbits, which includes HAT-P-11 b (Sanchis-Ojeda & Winn, 2011), GJ 436 b (Bourrier et al., 2018, 2022), WASP-107 b (Dai & Winn, 2017; Rubenzahl et al., 2021), and GJ 3470 b (Stefansson et al., 2022).
  • A possible explanation for the polar and eccentric (potentially eccentric in the case of HATS-38 b) orbits of these planets is high-eccentricity tidal migration produced by secular interactions from yet undetected outer planetary companions or the observed wide-orbit stars. If HATS-38b has a circular orbit, the polar orbit could also be explained with formation in a primordially misaligned disk. In contrast, we disfavor in-situ excitation of inclinations and eccentricities for either system, as this would require companions at orbital distances we have ruled out with out-of-transit RVs.
@@ Acknowledgments
We would like to thank Coel Hellier for the TRAPPIST and EulerCAM light curves of WASP-139. We thank Jiayin Dong for her help with the implementation of the Bayesian framework to derive the true 3D obliquity distributions. We also thank the anonymous referee for their thoughtful comments and suggestions that strengthened the manuscript. J.I.E.R and C.P. acknowledge support from ANID BASAL project FB210003. J.I.E.R. acknowledges support from ANID Doctorado Nacional grant 2021-21212378. C.P. acknowledges support from FONDECYT Project 1210425, CASSACA grant CCJRF2105, and ANID+REC Convocatoria Nacional subvención a la instalación en la Academia convocatoria 2020 PAI77200076. A.J. and R.B. acknowledge support from ANID – Millennium Science Initiative – ICN12_009 and AIM23-0001. R.B. acknowledges support from FONDECYT Project 1241963. A.J. acknowledges support from FONDECYT project 1210718. J.P.L. acknowledges co-funding by the European Union (ERC, FIERCE, 101052347), by FCT through national funds, and by FEDER through COMPETE2020 - Programa Operacional Competitividade e Internacionalização grants UIDB/04434/2020 and UIDP/04434/2020. M.T.P. acknowledges support from the ANID Post-doctoral fellowship 3210253.